OpenStudy (anonymous):

In ABC, if mA = 100° and mB = 50°, what is mC

OpenStudy (lucaz):

for any triangle, the sum of the interior angels is always 180 degrees A+B+C=180 100+50+C=180 C=180-150=30º

OpenStudy (anonymous):

Given ABC, with mB = 37° and mC = 49°, which of the following descriptions of the triangle are true